# Sequential giant planet formation initiated by disc substructure

### v1

@@ Abstract
Context. Planet formation models are necessary to understand the origins of diverse planetary systems. Circumstellar disc substructures have been proposed as preferred locations of planet formation but a complete formation scenario has not been covered by a single model so far.
@@
Methods. We connect dust coagulation and drift, planetesimal formation, $N$-body gravity, pebble accretion, planet migration, planetary gas accretion and gap opening in one consistent modelling framework.
Results. We find rapid formation of multiple gas giants from the initial disc substructure. The migration trap near the substructure allows the formation of cold gas giants. A new pressure maximum is created at the outer edge of the planetary gap, which triggers the next generation of planet formation resulting in a compact chain of giant planets. A high planet formation efficiency is achieved as the first gas giants are effective in preventing dust from drifting further inwards, which preserves materials for planet formation.
@@ 1 Introduction
Current planet formation models in general fail to meet both the physical and cosmochemical constraints required to explain the Solar System (e.g. Matsumura et al., 2017; Liu et al., 2019; Bitsch et al., 2019; Lau et al., 2024). The main challenge remains to be the ‘migration problem’, where rapid migration occurs for planetary cores of 1 to 10 $M_{\oplus}$ resulting in the formation of super-Earths and hot Jupiters. While the above models generally assume a smooth planetary disc, multiple works (e.g. Coleman & Nelson, 2016; Morbidelli, 2020; Guilera et al., 2020; Chambers, 2021; Andama et al., 2022; Lau et al., 2022) modelled the formation and evolution of planetary cores retained at the migration trap near a pressure bump in the disc. Nonetheless, the origin of such pressure bumps remains uncertain, the proposed non-planetary causes include late-stage infall of material (Gupta et al., 2023), sublimation (Saito & Sirono, 2011), instabilities (Takahashi & Inutsuka, 2014; Flock et al., 2015; Dullemond & Penzlin, 2018) and the edge of the dead zone, where the magneto-rotational instability (MRI) is suppressed (Pinilla et al., 2016).
@@
More recently, the high-resolution interferometry observations by the Atacama Large Millimeter/submillimeter Array (ALMA) show that substructure are typical in protoplanetary discs. Multiple surveys (e.g. Andrews et al., 2018; Long et al., 2018; Cieza et al., 2021), have shown that most of the substructures are presented in the form of axisymmetric rings. While these observations are limited to large and bright discs, disc population synthesis and theoretical models (Toci et al., 2021; Zormpas et al., 2022) demonstrate that disc substructures may be common in unresolved discs as well.
@@
Chatterjee & Tan (2013) presented an analytic model demonstrating the ‘inside-out’ planet formation scenario, where planet formation starts at the outer edge of the MRI active zone around the star. Although this work focuses on explaining the tightly-packed chains of planets commonly seen in exoplanetary systems, the model suggests the possibility of planet formation being triggered by the planet formed in the previous generation. A similar idea was also proposed for the formation of Saturn after the completion of Jupiter (Kobayashi et al., 2012), in which the core of Saturn grows rapidly without significant inward drift in the pressure bump induced by the planetary gap of Jupiter, although they did not consider planetesimal formation from dust and pebble accretion.
@@
Motivated by the current models and observations, substructure in protoplanetary discs has recently emerged as an ideal location for planet formation (Chambers, 2021; Lau et al., 2022; Jiang & Ormel, 2023). Lau et al. (2022) modelled the formation and evolution of planetesimals in an initial axisymmetric disc substructure by coupling the dust and gas evolution code DustPy (Stammler & Birnstiel, 2022) with the parallelized symplectic $N$-body integrator SyMBA parallelized (SyMBAp) (Lau & Lee, 2023). As the disc evolves and satisfies the condition for the streaming instability and the subsequent gravitational collapse, a fraction of the dust is converted into planetesimals as $N$-body particles. On top of the full $N$-body gravitational interactions, additional subroutines are gas drag, planet-disc interactions, and pebble accretion. Lau et al. (2022) showed the rapid formation of planetary cores thanks to the concentrated dust, which are also retained due to the migration trap near the pressure bump. However, this work did not attempt to form giant planets as they focused on the formation of planetary cores, where gas accretion and planetary gap opening are missing in the model.
@@
In the presented work, we further developed the model in Lau et al. (2022) for the formation of giant planets initiated by a disc substructure and found a scenario of sequential planet formation. In the following, Sec. 2 summarizes the methods adopted in Lau et al. (2022) and the new components implemented in this work. The results are presented in Sect. 3, which are followed by the discussions in Sect. 4. The findings of this work are summarized in Sect. 5.
@@ 2 Method
We employed the dust and gas evolution code DustPy v1.0.3 (Stammler & Birnstiel, 2022) and the symplectic $N$-body integrator SyMBAp v1.6 (Lau & Lee, 2023), a parallelized version of the Symplectic Massive Body Algorithm (SyMBA) (Duncan et al., 1998). The coupling of the two codes to construct a consistent planet formation model was first presented in Lau et al. (2022), which only modelled the formation of planetary cores. In this work, we added gas accretion and planetary gap opening to model the subsequent evolution of the embryos formed at a pressure bump. The following summarizes the employed method in Lau et al. (2022) and describes the new components in details.
@@ 2.1 Disc model
DustPy simulates the viscous evolution of the gas, coagulation, fragmentation, advection, and diffusion of the dust in a protoplanetary disc.
@@ 2.1.1 Gas component
$\Sigma_{\mathrm{g,init}}=\frac{M_{\mathrm{disc}}}{2\pi r_{\mathrm{c}}^{2}}\left(\frac{r}{r_{\mathrm{c}}}\right)^{-1}\exp\left(-\frac{r}{r_{\mathrm{c}}}\right)$ (1)
with the distance from the star $r$, the initial mass of the disc $M_{\mathrm{disc}}$ and the characteristic radius $r_{\mathrm{c}}$. We set $M_{\mathrm{disc}}=0.0263M_{\odot}$ and $r_{\mathrm{c}}=50\ \mathrm{au}$, which imply $\Sigma_{\mathrm{g,init}}\ (r=5\ \mathrm{au})\approx 134.6\ \mathrm{g}\ \mathrm{cm}^{-2}$.
@@
(Lüst, 1952; Lynden-Bell & Pringle, 1974), with the back reaction from the dust is neglected. The Shakura & Sunyaev (1973) $\alpha$-parametrization is adopted for the kinematic viscosity $\nu$ such that
$\nu=\alpha c_{\mathrm{s}}H_{\mathrm{g}}$ (3)
with the speed of sound $c_{\mathrm{s}}$ and the disc scale height $H_{\mathrm{g}}$. The viscosity parameter $\alpha=\{3,5\}\times 10^{-4}$ is set in this work. The disc scale height is defined by $H_{\mathrm{g}}\equiv c_{\mathrm{s}}/\Omega_{\mathrm{K}}$, where the local Keplerian orbital frequency $\Omega_{\mathrm{K}}=\sqrt{GM_{\odot}/r^{3}}$ with the gravitational constant $G$. The isothermal sound speed is used and given by $c_{\mathrm{s}}=\sqrt{k_{\mathrm{B}}T/\mu}$ with the Boltzmann constant $k_{\mathrm{B}}$, the midplane temperature $T$ and the mean molecular weight of the gas $\mu=2.3m_{\mathrm{p}}$. The disc is assumed to be passively irradiated by the Solar luminosity at a constant angle of 0.05, which gives a midplane temperature profile
@@
We note that the normalization is about 0.84 of that in Lau et al. (2022), which is the result of a correction made to DustPy since v1.0.2. This setup yields the dimensionless gas disc scale height
@@
$\eta=-\frac{\hat{h}_{\mathrm{g}}^{2}}{2}\frac{\partial\ln P}{\partial\ln r}$ (6)
with the midplane gas pressure $P$, which describes the degree of ‘sub-Keplerity’ of the gas. A logarithmic radial grid is adopted with with 133 cells from 3 to 53 au and with additional 42 cells from 53 to 1000 au.
@@ 2.1.2 Dust component
with the global dust-to-gas ratio $Z$ set at the Solar metallicity of 0.01. We follow Mathis et al. (1977), i.e. the MRN size distribution of the interstellar medium, for the initial size distribution of the dust grains. The maximum initial size is set at $1\ \mu\mathrm{m}$ with the internal density of $1.67\ \mathrm{g}\ \mathrm{cm}^{-3}$ assumed. A total of 141 dust mass bins logarithmically spaced from $10^{-12}$ to $10^{8}$ g are used. Each dust species is evolved in time through transport with the advection-diffusion equation (Clarke & Pringle, 1988) coupled to growth and fragmentation with the Smoluchowski equation. The fragmentation velocity is assumed to be $5\ \mathrm{m}\mathrm{s}^{-1}$. At collision velocities above which, the dust aggregates are assumed to fragment. The Stokes number Sti of the dust in each dust species $i$ is calculated by considering the Epstein and the Stokes I regimes. The dust scale height of each dust species $H_{\mathrm{d},i}$ is calculated according to Dubrulle et al. (1995)
$H_{\mathrm{d},i}=H_{\mathrm{g}}\sqrt{\frac{\alpha}{\alpha+\mathrm{St}_{i}}}$ (8)
@@ 2.1.3 Initial disc gap
An initial axisymmetric gap is introduced to the disc following the model by Dullemond et al. (2018), which is motivated by the commonly observed substructures in protoplanetary discs. To modify the gas profile, we applied a modified $\alpha$-parameter with radial dependence $\alpha^{\prime}(r)=\alpha/F(r)$, where the function
@@
Since we do not study the physical cause of the initial disc gap, the modified $\alpha^{\prime}(r)$-parameter only serves the purpose to attain a target gas profile and does not change the actual turbulence in the disc. Therefore, the modified $\alpha$-parameter $\alpha^{\prime}(r)$ is exclusively experienced by the gas, while dust diffusion, dust scale height and turbulent collision speeds are set by $\alpha$. Nonetheless, the dust evolves according to the resulting gas profile. We note that this treatment is not consistent with the substructure formation scenarios where the actual turbulence is changed, e.g. the edge of the dead-zone, but consistent with the cases where the turbulence is unchanged, e.g. infall of material.
@@ 2.1.4 Planetesimal formation
Figure 1: Differential mass distribution of the planetesimals formed near the initial disc substructure for the set of simulations with $\alpha=5\times 10^{-4}$ drawn from the adopted initial mass function by Gerbig & Li (2023). There are 10 bins per decade in mass $m$ and $\Delta N$ is the number of planetesimals in each bin. Each colour corresponds to one of the five simulations.
Lau et al. (2022) adopted the truncated power-law cumulative mass distribution from the fitting by Abod et al. (2019). However, the upper end of the distribution is not limited and Lau et al. (2024) noted that the largest planetesimal in the actual realization depends on the total number of planetesimals. Therefore, here we adopt the Toomre-like instability criterion $Q_{\mathrm{p}}$ for the gravitational collapse of the dense filament induced by streaming instability and the initial mass function from Gerbig & Li (2023). The criterion for collapse is $Q_{\mathrm{p}}<1$ with
@@
and the mass-averaged Stokes number of the dust in the cell $\mathrm{St}_{\mathrm{avg}}$. The local dust surface density $\Sigma_{\mathrm{d,local}}$ is assumed to be 10 times of the averaged $\Sigma_{\mathrm{d}}$ of the cell. Motivated by the streaming instability simulations in Schreiber & Klahr (2018), the small-scale diffusion parameter $\delta$ is set at $10^{-5}$. The model converts dust into planetesimals based on the prescription by Dra̧żkowska et al. (2016) and Schoonenberg et al. (2018). The $Q_{\mathrm{p}}$ criterion is combined with the smooth planetesimal formation activation function from Miller et al. (2021), which is given by
$\mathcal{P_{\mathrm{pf}}}=\frac{1}{1+\exp{[10\times(Q_{\mathrm{p}}-0.75)]}}$ (11)
@@
Then, the removed dust is summed over all dust species and added to the local planetesimal mass surface density. We first draw the location of a new planetesimal using the radial profile of the planetesimal mass surface density. Then, we draw the planetesimal mass according to the initial mass function given by Gerbig & Li (2023), which is resulting from the stability analysis of the dispersion relation for dust influenced by turbulent diffusion (Klahr & Schreiber, 2021). The maximum and minimum masses associated with the unstable modes are given by
$m_{\max,\min}=\frac{9}{8}\sqrt{\frac{\pi}{2}}\left(\frac{\delta}{\mathrm{St}_{\mathrm{avg}}}\right)^{3/2}\hat{h}_{\mathrm{g}}^{3}\left(\frac{1}{Q_{\mathrm{p}}}\pm\sqrt{\frac{1}{Q_{\mathrm{p}}^{2}}-1}\right)^{2}M_{\odot}$ (13)
@@
$\approx 0.37\left(\frac{\delta/10^{-5}}{\mathrm{St}_{\mathrm{avg}}/0.1}\right)^{3/2}\left(\frac{\hat{h}_{\mathrm{g}}}{0.05}\right)^{3}\left(\frac{Q_{\mathrm{p}}}{1}\right)^{2}M_{\mathrm{Ceres}}$ (15)
assuming isotropy in the small-scale diffusion. We refer the readers to Eq. (20) in Gerbig & Li (2023) for the expression of the probability density function. Figure 1 shows an example of the differential mass distribution of the planetesimals formed at about 6.5 au near the initial disc substructure for the set of simulations with $\alpha=5\times 10^{-4}$.
As described in Lau et al. (2022), the eccentricity $e$ and the inclination $i$ in radian are randomly drawn from two Rayleigh distributions with the scale parameters $10^{-6}$ and $5\times 10^{-7}$ respectively. The rest of the angles of the orbital elements in radian are drawn randomly from 0 to $2\pi$. The physical radius $R_{\mathrm{p}}$ is calculated by assuming an internal density $\rho_{\mathrm{s}}$ of 1.5 $\mathrm{g}\ \mathrm{cm}^{-3}$. The drawn mass is subtracted from the surface density from the nearest radial grid cells and the realization stops when the total remaining mass is less than the drawn mass. Any residue of the planetesimal mass surface density and the last drawn value of planetesimal mass is retained for the next time step to avoid bias towards the lower mass, i.e. the drawn mass will be realized as soon as enough planetesimal surface density is available.

### v2

@@ Abstract
Context. Planet formation models are necessary to understand the origins of diverse planetary systems. Circumstellar disc substructures have been proposed as preferred locations of planet formation, but a complete formation scenario has not been covered by a single model so far.
@@
Methods. We connect dust coagulation and drift, planetesimal formation, $N$-body gravity, pebble accretion, planet migration, planetary gas accretion, and gap opening in one consistent modelling framework.
Results. We find rapid formation of multiple gas giants from the initial disc substructure. The migration trap near the substructure allows for the formation of cold gas giants. A new pressure maximum is created at the outer edge of the planetary gap, which triggers the next generation of planet formation resulting in a compact chain of giant planets. A high planet formation efficiency is achieved, as the first gas giants are effective at preventing dust from drifting further inwards, which preserves material for planet formation.
@@ 1 Introduction
Current planet formation models in general fail to meet both the physical and cosmochemical constraints required to explain the Solar System (e.g. Matsumura et al., 2017; Liu et al., 2019; Bitsch et al., 2019; Lau et al., 2024). The main challenge remains to be the ‘migration problem’, where rapid migration occurs for planetary cores of 1 to 10 $M_{\oplus}$, resulting in the formation of super-Earths and hot Jupiters. While the above models generally assume a smooth planetary disc, multiple works (e.g. Coleman & Nelson, 2016; Morbidelli, 2020; Guilera et al., 2020; Chambers, 2021; Andama et al., 2022; Lau et al., 2022) have modelled the formation and evolution of planetary cores retained at the migration trap near a pressure bump in the disc. Nonetheless, the origin of such pressure bumps remains uncertain, the proposed non-planetary causes include late-stage infall of material (Gupta et al., 2023), sublimation (Saito & Sirono, 2011), instabilities (Takahashi & Inutsuka, 2014; Flock et al., 2015; Dullemond & Penzlin, 2018), and the edge of the dead zone, where the magneto-rotational instability (MRI) is suppressed (Pinilla et al., 2016).
@@
More recently, the high-resolution interferometry observations by the Atacama Large Millimeter/submillimeter Array (ALMA) have shown that substructures are typical in protoplanetary discs. Multiple surveys (e.g. Andrews et al., 2018; Long et al., 2018; Cieza et al., 2021), have shown that most of the substructures are presented in the form of axisymmetric rings. While these observations are limited to large and bright discs, disc population synthesis and theoretical models (Toci et al., 2021; Zormpas et al., 2022) have demonstrated that disc substructures may be common in unresolved discs as well.
@@
Chatterjee and Tan (2013) presented an analytic model demonstrating the ‘inside-out’ planet formation scenario, where planet formation starts at the outer edge of the MRI active zone around the star. Although this work focuses on explaining the tightly packed chains of planets commonly seen in exoplanetary systems, the model suggests the possibility of planet formation being triggered by the planet formed in the previous generation. A similar idea was also proposed for the formation of Saturn after the completion of Jupiter (Kobayashi et al., 2012), in which the core of Saturn grows rapidly without significant inward drift in the pressure bump induced by the planetary gap of Jupiter, although they did not consider planetesimal formation from dust and pebble accretion.
@@
Motivated by the current models and observations, a substructure in a protoplanetary disc has recently emerged as an ideal location for planet formation (Chambers, 2021; Lau et al., 2022; Jiang & Ormel, 2023). Lau et al. (2022) modelled the formation and evolution of planetesimals in an initial axisymmetric disc substructure by coupling the dust and gas evolution code DustPy (Stammler & Birnstiel, 2022) with the parallelised symplectic $N$-body integrator SyMBA parallelised (SyMBAp; Lau & Lee, 2023). As the disc evolves and satisfies the condition for the streaming instability and the subsequent gravitational collapse, a fraction of the dust is converted into planetesimals as $N$-body particles. On top of the full $N$-body gravitational interactions, additional subroutines are gas drag, planet-disc interactions, and pebble accretion. Lau et al. (2022) showed the rapid formation of planetary cores thanks to the concentrated dust, which are also retained due to the migration trap near the pressure bump. However, this work did not attempt to form giant planets, as the authors focused on the formation of planetary cores, where gas accretion and planetary gap opening are missing in the model.
@@
In this work, we further developed the model in Lau et al. (2022) for the formation of giant planets initiated by a disc substructure and found a scenario of sequential planet formation. In the following, Sec. 2 summarises the methods adopted in Lau et al. (2022) and the new components implemented in this work. The results are presented in Sect. 3, which are followed by the discussions in Sect. 4. The findings of this work are summarised in Sect. 5.
@@ 2 Method
We employed the dust and gas evolution code DustPy v1.0.3 (Stammler & Birnstiel, 2022) and the symplectic $N$-body integrator SyMBAp v1.6 (Lau & Lee, 2023), a parallelised version of the Symplectic Massive Body Algorithm (SyMBA; Duncan et al., 1998). The coupling of the two codes to construct a consistent planet formation model was first presented in Lau et al. (2022), which only modelled the formation of planetary cores. In this work, we added gas accretion and planetary gap opening to model the subsequent evolution of the embryos formed at a pressure bump. The following summarises the employed method in Lau et al. (2022) and describes the new components in detail.
@@ 2.1 Disc model
DustPy simulates the viscous evolution of the gas, coagulation, fragmentation, advection, and diffusion of the dust in a protoplanetary disc. The different parts of the disc model are described in this section.
@@ 2.1.1 Gas component
$\Sigma_{\mathrm{g,init}}=\frac{M_{\mathrm{disc}}}{2\pi r_{\mathrm{c}}^{2}}\left(\frac{r}{r_{\mathrm{c}}}\right)^{-1}\exp\left(-\frac{r}{r_{\mathrm{c}}}\right),$ (1)
with the distance from the star $r$, the initial mass of the disc $M_{\mathrm{disc}}$, and the characteristic radius $r_{\mathrm{c}}$. We set $M_{\mathrm{disc}}=0.0263M_{\odot}$ and $r_{\mathrm{c}}=50\ \mathrm{au}$, which imply $\Sigma_{\mathrm{g,init}}\ (r=5\ \mathrm{au})\approx 134.6\ \mathrm{g}\ \mathrm{cm}^{-2}$.
@@
(Lüst, 1952; Lynden-Bell & Pringle, 1974), while the backreaction from the dust is neglected. The Shakura and Sunyaev (1973) $\alpha$-parametrisation was adopted for the kinematic viscosity $\nu$ such that
$\nu=\alpha c_{\mathrm{s}}H_{\mathrm{g}},$ (3)
with the speed of sound $c_{\mathrm{s}}$ and the disc scale height $H_{\mathrm{g}}$. The viscosity parameter $\alpha=\{3,5\}\times 10^{-4}$ was set in this work. The disc scale height is defined by $H_{\mathrm{g}}\equiv c_{\mathrm{s}}/\Omega_{\mathrm{K}}$, where the local Keplerian orbital frequency $\Omega_{\mathrm{K}}=\sqrt{GM_{\odot}/r^{3}}$ with the gravitational constant $G$. The isothermal sound speed was used and given by $c_{\mathrm{s}}=\sqrt{k_{\mathrm{B}}T/\mu}$ with the Boltzmann constant $k_{\mathrm{B}}$, the midplane temperature $T$ and the mean molecular weight of the gas $\mu=2.3m_{\mathrm{p}}$. The disc was assumed to be passively irradiated by the solar luminosity at a constant angle of 0.05, which gives a midplane temperature profile
@@
We note that the normalisation is about 0.84 of that in Lau et al. (2022), which is the result of a correction made to DustPy since v1.0.2. This setup yields the dimensionless gas disc scale height
@@
$\eta=-\frac{\hat{h}_{\mathrm{g}}^{2}}{2}\frac{\partial\ln P}{\partial\ln r},$ (6)
with the midplane gas pressure $P$, which describes the degree of ‘sub-Keplerity’ of the gas. A logarithmic radial grid was adopted with with 133 cells from 3 to 53 au and with an additional 42 cells from 53 to 1000 au.
@@ 2.1.2 Dust component
with the global dust-to-gas ratio $Z$ set at the solar metallicity of 0.01. We followed Mathis et al. (1977), that is, the MRN size distribution of the interstellar medium, for the initial size distribution of the dust grains. The maximum initial size was set at $1\ \mu\mathrm{m}$ with the internal density of $1.67\ \mathrm{g}\ \mathrm{cm}^{-3}$ assumed. A total of 141 dust mass bins logarithmically spaced from $10^{-12}$ to $10^{8}$ g were used. Each dust species was evolved in time through transport with the advection-diffusion equation (Clarke & Pringle, 1988) coupled to growth and fragmentation with the Smoluchowski equation. The fragmentation velocity was assumed to be $5\ \mathrm{m}\mathrm{s}^{-1}$. At collision velocities above which, the dust aggregates are assumed to fragment. The Stokes number Sti of the dust in each dust species $i$ was calculated by considering the Epstein and the Stokes I regimes. The dust scale height of each dust species $H_{\mathrm{d},i}$ was calculated according to Dubrulle et al. (1995),
$H_{\mathrm{d},i}=H_{\mathrm{g}}\sqrt{\frac{\alpha}{\alpha+\mathrm{St}_{i}}},$ (8)
@@ 2.1.3 Initial disc gap
An initial axisymmetric gap was introduced to the disc following the model by Dullemond et al. (2018), which is motivated by the commonly observed substructures in protoplanetary discs. To modify the gas profile, we applied a modified $\alpha$-parameter with radial dependence $\alpha^{\prime}(r)=\alpha/F(r)$, where the function
@@
Since we do not study the physical cause of the initial disc gap, the modified $\alpha^{\prime}(r)$-parameter only serves the purpose to attain a target gas profile and does not change the actual turbulence in the disc. Therefore, the modified $\alpha$-parameter $\alpha^{\prime}(r)$ is exclusively experienced by the gas, while dust diffusion, dust scale height and turbulent collision speeds are set by $\alpha$. Nonetheless, the dust evolves according to the resulting gas profile. We note that this treatment is not consistent with the substructure formation scenarios where the actual turbulence is changed, for example, the edge of the dead-zone, but consistent with the cases where the turbulence is unchanged, for example, infall of material.
@@ 2.1.4 Planetesimal formation
Figure 1: Differential mass distribution of the planetesimals formed near the initial disc substructure for the set of simulations with $\alpha=5\times 10^{-4}$ drawn from the adopted initial mass function by Gerbig and Li (2023). There are 10 bins per decade in mass $m$ and $\Delta N$ is the number of planetesimals in each bin. Each colour corresponds to one of the five simulations.
Lau et al. (2022) adopted the truncated power-law cumulative mass distribution from the fitting by Abod et al. (2019). However, the upper end of the distribution is not limited and Lau et al. (2024) noted that the largest planetesimal in the actual realisation depends on the total number of planetesimals. Therefore, here we adopted the Toomre-like instability criterion $Q_{\mathrm{p}}$ for the gravitational collapse of the dense filament induced by streaming instability and the initial mass function from Gerbig and Li (2023). The criterion for collapse is $Q_{\mathrm{p}}<1$ with
@@
and the mass-averaged Stokes number of the dust in the cell $\mathrm{St}_{\mathrm{avg}}$. The local dust surface density $\Sigma_{\mathrm{d,local}}$ was assumed to be 10 times of the averaged $\Sigma_{\mathrm{d}}$ of the cell. Motivated by the streaming instability simulations in Schreiber & Klahr (2018), the small-scale diffusion parameter $\delta$ was set at $10^{-5}$. The model converts dust into planetesimals based on the prescription by Dra̧żkowska et al. (2016) and Schoonenberg et al. (2018). The $Q_{\mathrm{p}}$ criterion was combined with the smooth planetesimal formation activation function from Miller et al. (2021), which is given by
$\mathcal{P_{\mathrm{pf}}}=\frac{1}{1+\exp{[10\times(Q_{\mathrm{p}}-0.75)]}},$ (11)
@@
Then, the removed dust is summed over all dust species and added to the local planetesimal mass surface density. We first draw the location of a new planetesimal using the radial profile of the planetesimal mass surface density. Then, we draw the planetesimal mass according to the initial mass function given by Gerbig and Li (2023), which is resulting from the stability analysis of the dispersion relation for dust influenced by turbulent diffusion (Klahr & Schreiber, 2021). The maximum and minimum masses associated with the unstable modes are given by
$m_{\max,\min}=\frac{9}{8}\sqrt{\frac{\pi}{2}}\left(\frac{\delta}{\mathrm{St}_{\mathrm{avg}}}\right)^{3/2}\hat{h}_{\mathrm{g}}^{3}\left(\frac{1}{Q_{\mathrm{p}}}\pm\sqrt{\frac{1}{Q_{\mathrm{p}}^{2}}-1}\right)^{2}M_{\odot},$ (13)
@@
$\approx 0.37\left(\frac{\delta/10^{-5}}{\mathrm{St}_{\mathrm{avg}}/0.1}\right)^{3/2}\left(\frac{\hat{h}_{\mathrm{g}}}{0.05}\right)^{3}\left(\frac{Q_{\mathrm{p}}}{1}\right)^{2}M_{\mathrm{Ceres}},$ (15)
assuming isotropy in the small-scale diffusion. We refer the readers to Eq. (20) in Gerbig and Li (2023) for the expression of the probability density function. Figure 1 shows an example of the differential mass distribution of the planetesimals formed at about 6.5 au near the initial disc substructure for the set of simulations with $\alpha=5\times 10^{-4}$.
As described in Lau et al. (2022), the eccentricity $e$ and the inclination $i$ in radian are randomly drawn from two Rayleigh distributions with the scale parameters $10^{-6}$ and $5\times 10^{-7}$ respectively. The rest of the angles of the orbital elements in radian are drawn randomly from 0 to $2\pi$. The physical radius $R_{\mathrm{p}}$ is calculated by assuming an internal density $\rho_{\mathrm{s}}$ of 1.5 $\mathrm{g}\ \mathrm{cm}^{-3}$. The drawn mass is subtracted from the surface density from the nearest radial grid cells and the realisation stops when the total remaining mass is less than the drawn mass. Any residue of the planetesimal mass surface density and the last drawn value of planetesimal mass is retained for the next time step to avoid bias towards the lower mass, that is, the drawn mass will be realised as soon as enough planetesimal surface density is available.